OpenStudy (anonymous):

Can a spherical tank with a 50ft diameter hold enough water to fill a swimming pool that has a 50m length, a 25m width, and a constant 2m depth? Explain your answer and provide work and calculations to support your response

OpenStudy (anonymous):

d = 50 feet = 15,24 m => r = 7,62 m V(tank/sphere) = (4pi7,62^2)/3 = 1853,3 m^3 V(pool/rectangular solid) = 50*25*2 = 2500 m^3 V(tank)<V(pool), so, the tank would not fill the pool...

OpenStudy (anonymous):

Thanks so much for your help!

OpenStudy (anonymous):

No problem. I noticed I made a typo, though. Sphere volume is (4*pi*7,62^3)/3. not to the power of two.